THE ARGON-CYCLOPENTADIENYL THALLIUM WEAKLY-BOUND COMPLEX, ROTATIONAL SPEC-TRUM AND STRUCTURE

CHAKREE TANJAROON, ADAM DALY, and \underline{STEPHEN G. KUKOLICH}, Department of Chemistry, The University of Arizona, Tucson, Arizona 85721..

The symmetric-top rotational spectrum of the neutral argon-cyclopentadienyl thallium complex was measured using a Flygare-Balle-type Fourier transform microwave spectrometer. This appears to be the first rotational spectrum for a noble gas-organometallic complex. The rotational constants are B = 372.45 MHz for Ar-C_{5}H_{5}^{205}Tl and 373.35 MHz for Ar-C_{5}H_{5}^{203}Tl isotopomers. D_{J} = 0.1 kHz and D_{JK} = 0.4 kHz, indicating a fairly rigid structure. The argon atom is located on top of the cyclopentadienyl ring, on the a-axis of the C_{5}H_{5}Tl monomer, opposite the thallium atom. The separation distance between argon and the cyclopentadienyl ring is 3.56 Å. The binding energy of Ar-C_{5}H_{5}Tl was estimated to be about 1-3 kJ mol^{−1} using low-level MP2 theory.